In a study of 287 comma 689 cell phone​ users, it was found that 31 developed cancer of the brain or nervous system. Assuming that cell phones have no​ effect, there is a 0.000115 probability of a person developing cancer of the brain or nervous system. We therefore expect about 34 cases of such cancer in a group of 287 comma 689 people. Estimate the probability of 31 or fewer cases of such cancer in a group of 287 comma 689 people. What do these results suggest about media reports that cell phones cause cancer of the brain or nervous​ system?
Answer #1
n= 287689 p= 0.000115
here mean of distribution=μ=np= 33.0842
and standard deviation σ=sqrt(np(1-p))= 5.7516
for normal distribution z score =(X-μ)/σx
therefore from normal approximation of binomial distribution:

probability of 31 or fewer cases :

probability = P(X<31) = P(Z<-0.36)= 0.3594

(please try 0.3897 or 0.3585 if above comes wrong and revert)